An incandescent lightbulb has a tungsten filament that is heated to a temperature of 3.36 x 103 K when an electric current passes through it. If the surface area of the filament is approximately 0.519 x 10−4 m2 and it has an emissivity of 0.320, what is the power radiated by the bulb?
Answer #1

We use the power radiated formula:

H = A\sigma \varepsilon T^4

using given data:

H =0.519\times 10^{-4}(5.67\times 10^{-8})(0.320)(3.36\times 10^3)^4 = 120.02 W